OpenStudy (anonymous):

solve for all values of x \[1/4^{1-2x} =32\]

OpenStudy (zarkon):

\[\frac{1}{4^{1-2x}} =32\] \[\frac{1}{2^{2(1-2x)}} =2^5\] \[2^{-2(1-2x)} =2^5\] \[-2(1-2x) =5\] \[1-2x =-\frac{5}{2}\] \[-2x =-\frac{5}{2}-1=-\frac{7}{2}\] \[x=\frac{7}{4}\]

OpenStudy (anonymous):

thank you very much :)